Does the general quintic equation (for instance Bring's x5 + px + q = 0) have roots?
Answer by venugopalramana(3286)

Does the general quintic equation (for instance Bring's x5 + px + q = 0) have roots?
WHY NOT ?WE CAN MAKE ANY EQN.WITH THE ROOTS WE WANT
TAKE YOUE OWN EXAMPLE
F(X)=X^5+PX+Q=0
IF WE WANT TO MAKE X=1 A ROOT ,WE PROCEED AS FOLLOWS...FIND F(1)
F(1)=1+P+Q...SO P+AQ=0...SO WE CAN TAKE...P=-Q AND THEN IT WILL HAVE A ROOT OF X=1.
IS THIS YOUR QUESTION?OR YOU HAVE SOME THING ELSE ON YOUR MIND?